# Social Bias in Popular Question-Answering Benchmarks

### v2

@@ Abstract
Question-answering (QA) and reading comprehension (RC) benchmarks are essential for assessing the capabilities of large language models (LLMs) in retrieving and reproducing knowledge. However, we demonstrate that popular QA and RC benchmarks are biased and do not cover questions about different demographics or regions in a representative way, potentially due to a lack of diversity of those involved in their creation. We perform a qualitative content analysis of 30 benchmark papers and a quantitative analysis of 20 respective benchmark datasets to learn (1) who is involved in the benchmark creation, (2) how social bias is addressed or prevented, and (3) whether the demographics of the creators and annotators correspond to particular biases in the content. Most analyzed benchmark papers provided insufficient information regarding the stakeholders involved in benchmark creation, particularly the annotators. Notably, just one of the benchmark papers explicitly reported measures taken to address social representation issues. Moreover, the data analysis revealed gender, religion, and geographic biases across a wide range of encyclopedic, commonsense, and scholarly benchmarks. More transparent and bias-aware QA and RC benchmark creation practices are needed to facilitate better scrutiny and incentivize the development of fairer LLMs.
@@
Angelie Kraft University of Hamburg Leuphana University Lüneburg Weizenbaum Institute angelie.kraft@leuphana.de                        Judith Simon University of Hamburg                        Sonja Schimmler TU Berlin Fraunhofer FOKUS Weizenbaum Institute
@@ 1 Introduction
Large language models (LLMs) inhabit the core of a wide range of user-facing systems, powering applications such as chatbots, which are utilized as writing and coding assistants, search engines, and advisors. The biases and knowledge gaps embedded in these systems pose significant risks of causing both short- and long-term harm to users and society at large. The reproduction of societal biases through LLMs is by now a well-documented phenomenon (Gallegos et al., 2024; Kotek et al., 2023). Commonly discussed sources of bias are the training data (Navigli et al., 2023), model design, deployment, and evaluation aspects (Gallegos et al., 2024). Indeed, optimizing LLMs to perform well on popular benchmarks is highly incentivized, as strong performance can enhance a researcher’s visibility and credibility (Koch et al., 2021). However, if these widely used benchmarks are biased, they effectively incentivize model optimization towards biased standards (Bowman and Dahl, 2021). Our work provides one of the first systematic analyses demonstrating that many of the most widely adopted LLM benchmarks are, in fact, quietly unrepresentative.
Raji et al. (2021, p. 2) "describe a benchmark as a particular combination of a dataset or sets of datasets […], and a metric, conceptualized as representing one or more specific tasks or sets of abilities." They argue that popular benchmarks – while claiming universality – are actually limited in coverage and validity. Citing Haraway (2016), they criticize that this universality claim masks an inevitable positionality and value-ladenness which manifests itself in the lack of coverage of "non-Western contexts," under-representation of non-cis gender identities, and non-white racial identities. Yet, comprehensive empirical analyses of benchmark biases are sparse and mostly limited to benchmarks that are themselves dedicated to the measurement of bias instead of downstream task performance (Powers et al., 2024; Demchak et al., 2024).
@@
Our work aims to fill this gap and focuses on question-answering (QA) and reading comprehension (RC) benchmarks, i.e., tasks where the model is presented an explicit question and its generated answer is then checked for correctness (e.g., open-ended, fill-in-the-gap, or multiple choice;  Rogers et al., 2023). We argue that these tasks are the most direct proxies to the ways in which users query chatbots to gather information and, thus, the ways in which LLMs are shaping modern knowledge ecosystems.
Building on the definition of an AI benchmark introduced by Raji et al. (2021), we define a socially biased QA or RC benchmark as one that exhibits a statistical skew in the occurrence of demographic and/or geographic identifiers or names within its dataset, corresponding to societal gradients of power and injustice, such as the under-representation of non-cis-male gender identities or non-Western individuals, locations, or events. We argue that such biases in QA and RC benchmarks can cause societal harm. By overlooking marginalized demographics and geographies in evaluation, these benchmarks encourage the optimization of knowledge-driven language technologies to favor the interests of a privileged few, exacerbating injustices related to knowledge, i.e., epistemic injustice Fricker (2007). This injustice unfolds in two forms: hermeneutic injustice, where marginalized groups lack the knowledge resources to interpret their social experiences, and testimonial injustice, where their credibility is diminished due to prejudiced judgments (Fricker, 2007). Biased benchmarks not only widen gaps in collective knowledge resources but also privilege certain knowledges over others, designating them as more desirable for LLMs to reproduce.
Our work seeks to answer the following research questions:
@@
    Are potential social biases avoided or addressed in the creation of the benchmarks?
@@
    Are potential social biases in the datasets reflected in the demographics of the individuals involved in the benchmark creation process?
Based on a manual analysis of the 30 most popular QA and RC benchmark papers as well as a quantitative data analysis of 20 benchmark datasets, we identified (a) a lack of transparency regarding the individuals involved in benchmark dataset creation, (b) a lack of intentional prevention of biases, and (c) prominent gender, occupation, religion, and location biases for encyclopedic, commonsense, maths, and science benchmarks. In unveiling these issues, our work adds to the mounting criticism of current AI evaluation practices and shines a light on biased benchmarks being a potential source of LLM bias by incentivizing biased inference heuristics. [^1] The source code can be found here: https://github.com/krangelie/qa-benchmark-biases
@@ 2 Related Works
LLMs reproduce stereotypical associations (Nadeem et al., 2021; Kotek et al., 2023) and achieve different levels of accuracy for examples referring to different social groups in downstream-tasks (Park et al., 2018; Kiritchenko and Mohammad, 2018), such as QA (Parrish et al., 2022; Jin et al., 2024). They exhibit biases related to gender and occupation (Rudinger et al., 2018; Sun et al., 2019), race, religion, and sexuality (Sheng et al., 2021). These biases can lead to representational and allocational harms Barocas et al. (2017); Blodgett et al. (2020). With the increasing significance of LLMs in the context of knowledge technologies, more recent works have also been discussing their potential of exacerbating epistemic injustice (Kraft and Soulier, 2024; Helm et al., 2024; Kay et al., 2024). Sources of bias are non-representative training data, the training or inference algorithm, the deployment context and user interface, as well as evaluation with unrepresentative benchmarks (Gallegos et al., 2024; Suresh and Guttag, 2021; Bowman and Dahl, 2021). Age, gender, race, educational background, and first language of an annotator can influence their annotations and, consequently, the ground truths used to train and evaluate models (Pei and Jurgens, 2023; Al Kuwatly et al., 2020). Hence, crowdworker samples with low demographic diversity produce datasets of correspondingly low diversity and generalizability (Geva et al., 2019). Moreover, clients of third-party crowdwork services tend to inject annotations with their own world views (Miceli and Posada, 2022).
@@
Transparent documentation practices of datasets, including their biases and limitations, have been promoted as an important measure to prevent harmful outcomes (Bender and Friedman, 2018; Stoyanovich and Howe, 2019; Gebru et al., 2021), i.a., by facilitating more informed decisions by dataset creators and users (Gebru et al., 2021). Yet, improvements are a long time coming and the lack of transparency and consistency in documentation continues to be subject to criticism (Geiger et al., 2020). Reuel et al. (2024), recently proposed a structured AI benchmark assessment with criteria addressing aspects of design, implementation, documentation, maintenance, and retirement. They applied this schema to 24 foundation and non-foundation model benchmarks, covering natural language processing (NLP), agentic and ethical behavior benchmarks and found overall low levels of reproducibility and interpretability. MMLU scores lowest in their overall assessment. Our work sits in the same category but targets a social bias-related appraisal of benchmarks. Researchers in the area of algorithmic bias research have been investigating the biases of bias benchmarks, like BBQ (Powers et al., 2024; Parrish et al., 2022), BOLD and SAGED (Demchak et al., 2024; Dhamala et al., 2021; Guan et al., 2024). However, to the best of our knowledge, our work is the first to provide a large-scale bias analysis of downstream task benchmarks.
@@ 3.1 Benchmark Selection
To identify popular QA and RC benchmarks, we firstly selected all benchmarks including textual data (not excluding multimodal datasets) in the Papers with Code (PwC) corpus of machine learning dataset metadata [^2] https://paperswithcode.com/about, accessed: September 17, 2024 and ranked them by their citation counts. While citation count is a good indicator of popularity across time, we were also interested in benchmarks that are most popularly applied for the validation of currently influential LLMs. To identify such, we selected the most highly ranked models on the Chatbot Arena LLM Leaderboard (Chiang et al., 2024), [^3] https://lmarena.ai/, accessed: September 18, 2024 as well as the language models with the most likes on HuggingFace. [^4] https://huggingface.co/models?sort=likes, accessed: September 13, 2024 We extracted the top 20 models from both lists and collected all of the 40 related reports, i.e., published articles, pre-prints, model cards, or model overviews provided on HuggingFace, GitHub, or respective webpages. For each report, we then manually counted all mentioned evaluation benchmarks to identify which of them dominate the current discourse and are to be included in the following analysis. Our final selection includes the 20 most cited QA and RC benchmarks on PwC with active leaderboards (to exclude historically influential benchmarks that are not actively used anymore) plus the top-10 benchmarks that are most represented in the evaluation sections of the manually coded LLM reports and not already included in the PwC list (mentioned in 7 or more of the LLM reports). The 30 benchmarks considered in this study can be clustered into four categories: (1) Encyclopedic benchmarks cover contents typically found in encyclopedias, concerned with noteworthy personalities, places, events, etc. Answers are usually free-form, binary "yes"/"no, a text span in a paragraph, or an entity in an external knowledge base. (2) Commonsense benchmarks pose questions about everyday knowledge, e.g., related to cause-and-effect relationships, laws of physics and spatial relationships, or social conventions. Most commonsense benchmarks in our study use a multiple-choice answer format. (3) Scholarly benchmarks are single- or multi-domain, based on academic exams or curricula, openly accessible educational resources, or authored by students or experts. Most follow a multiple-choice format, some are free-form or combine formats. (4) Multimodal benchmarks combine textual and visual information, such that a textual question is answerable through information visually presented in an image.
@@ 3.2 Analysis of Benchmark Papers
Figure 1 gives a schematic overview of our benchmark paper analysis procedure. Guided by our research questions and following a content analysis approach similar to Birhane et al. (2022), we firstly coded all of the benchmark papers, i.e., research articles or pre-prints introducing the benchmark (as the main contribution, as a byproduct to a technical work, or as a test split to a new corpus). Using MAXQDA (VERBI Software, 2024), our first author firstly highlighted sections relevant to our research questions and suggested preliminary annotation labels on the fly. After the first phase of annotations, labels were merged and categorized where possible to create a codebook. This initial codebook was discussed in a workshop with four participants, including two of the authors and two colleagues from the TU Berlin. Discussions regarding the codebook content later informed its refinement and finalization. [^5] During the workshop, the codebook was presented as a list of labels with short descriptions and the external participants were asked to annotate two benchmark papers by marking and labeling text spans using this list. It, however, required a long time for the new annotators to comprehend the list of possible labels and understand the type of insights we were looking for. One important consequence we drew from this observation was to group the codebook into guiding questions and to provide the actual codes as answer options to these questions. This helped to accelerate the on-boarding. The final codebook was reformatted and implemented as an online questionnaire via LimeSurvey (LimeSurvey Project Team / Carsten Schmitz, 2012), which we used for the second wave of annotations. [^6] The full questionnaire is available in our repository. Using the final coding schema, all 30 benchmark papers were re-annotated by our first author and one external annotator each. We distributed the co-annotation among 12 experts, of which nine were PhD students with a research focus on NLP and QA, two were Master’s students and one was a medical professional–all with a working knowledge of NLP and experience in reading scientific texts. All annotators (including workshop participants and respective authors) were aged between 25 and 60, originating from India, Pakistan, China, Germany, and Kazakhstan. Roughly one third identified as female. [^7] All annotators (including workshop participants) were informed about the conditions and rights (including GDPR) upon participating in our study and all provided their written consent prior to participation. Their demographic details were collected in a separate questionnaire including a separate informed consent form.
@@ 3.3 Analysis of Benchmark Datasets
For the quantitative analysis of social bias within the benchmark datasets, we retrieved external information about entities (people, places, events, etc.) mentioned in the question-answer pairs from Wikidata, [^8] https://www.wikidata.org in particular, gender, occupation, religion, and location properties.
@@
Our analyses comprise two different scenarios depicted in Figure 2: Scenario 1: The questions or answers of benchmarks like NaturalQuestions and TriviaQA include entities described in Wikipedia articles and respective identifiers (e.g., article titles or URLs) are provided. Using these identifiers, we queried the Wikipedia API [^9] https://www.mediawiki.org/wiki/API:Main_page to retrieve the corresponding Wikidata QIDs. Using SPARQL, [^10] https://www.w3.org/TR/rdf-sparql-query/ we then retrieved properties of interest for these QIDs directly from the Wikidata knowledge graph, e.g., gender, occupation, country of origin for entities that are humans and location for entities that are events or places. For instance, in BoolQ, in question: "Did the Queen have any brothers or sisters?", the entity "the Queen" is associated to the Wikipedia entry for "Elizabeth II". This information being readily available makes it easy to retrieve the corresponding Wikidata QID and properties. Scenario 2: For benchmarks that are not by default linked to Wikipedia, we applied Entity Linking (EL) against Wikidata using the ReFinED model (Ayoola et al., 2022). [^11] We used the implementation available here: https://github.com/amazon-science/ReFinED (license: Apache 2.0). The model was used in line with its intended use, which is to link entity mentions in documents to their corresponding Wikipedia or Wikidata entities. This allowed us to identify entities mentioned in the questions and retrieve their respective Wikidata identifiers. We then followed the same steps as in the first scenario. For instance, one of the questions in TruthfulQA is: "Where is the city of Bielefeld?" [^12] A correct answer to this question is ”Bielefeld is in Germany” and an expected incorrect answer is ”Bielefeld does not exist”. The entity linker identifies "Bielefeld" as a Wikidata entity with the QID "Q2112", allowing to query properties associated with it from the Wikidata knowledge graph (e.g., its coordinates). We excluded BioASQ-QA, MATH, and the multimodal benchmarks from the analysis, as identifying social biases within these benchmarks would necessitate additional domain-specific expertise or extensive annotation efforts beyond the scope of this study. Some other benchmark datasets were excluded if they yielded less than 30 results for each property. A total of 20 benchmark datasets were included in our final quantitative analysis.

### v3

@@ Abstract
Question-answering (QA) and reading comprehension (RC) benchmarks are commonly used for assessing the capabilities of large language models (LLMs) to retrieve and reproduce knowledge. However, we demonstrate that popular QA and RC benchmarks do not cover questions about different demographics or regions in a representative way. We perform a content analysis of 30 benchmark papers and a quantitative analysis of 20 respective benchmark datasets to learn (1) who is involved in the benchmark creation, (2) whether the benchmarks exhibit social bias, or whether this is addressed or prevented, and (3) whether the demographics of the creators and annotators correspond to particular biases in the content. Most benchmark papers analyzed provide insufficient information about those involved in benchmark creation, particularly the annotators. Notably, just one (WinoGrande) explicitly reports measures taken to address social representation issues. Moreover, the data analysis revealed gender, religion, and geographic biases across a wide range of encyclopedic, commonsense, and scholarly benchmarks. Our work adds to the mounting criticism of AI evaluation practices and shines a light on biased benchmarks being a potential source of LLM bias by incentivizing biased inference heuristics.
@@
Angelie Kraft University of Hamburg Leuphana University Lüneburg Weizenbaum Institute angelie.kraft@uni-hamburg.de          Judith Simon University of Hamburg          Sonja Schimmler Technical University Berlin Fraunhofer FOKUS Weizenbaum Institute
@@ 1 Introduction
Large language models (LLMs) inhabit the core of a wide range of user-facing systems. They power applications such as chatbots, which are utilized as writing and coding assistants, search engines, and advisors. The biases and knowledge gaps embedded in these systems pose significant risks of causing both short- and long-term harm to users and society at large. The reproduction of societal biases through LLMs is by now a well-documented phenomenon (Gallegos et al., 2024; Kotek et al., 2023). Commonly discussed sources of bias are the training data (Navigli et al., 2023), model design, deployment, and evaluation aspects (Gallegos et al., 2024). Indeed, optimizing LLMs to perform well on popular benchmarks is highly incentivized, as strong performance can enhance a researcher’s visibility and credibility (Koch et al., 2021). However, it has been theorized that many widely used benchmarks are biased and effectively incentivize model optimization towards biased standards (Bowman and Dahl, 2021; Raji et al., 2021).
@@
Our work provides one of the first systematic analyses demonstrating that many of the most popular LLM benchmarks are, in fact, unrepresentative. Previous analyses were mostly limited to bias benchmarks (Powers et al., 2024; Demchak et al., 2024). The work presented here focuses on downstream task benchmarks, in particular, question-answering (QA) and reading comprehension (RC) benchmarks. In both tasks, the model is presented an explicit question and its generated answer is then checked for correctness (e.g., open-ended, fill-in-the-gap, or multiple choice;  Rogers et al., 2023). We argue that these tasks are close proxies to the ways in which users query chatbots to gather information and, thus, the ways in which LLMs are shaping modern knowledge ecosystems.
Raji et al. (2021, p. 2) "describe a benchmark as a particular combination of a dataset or sets of datasets […], and a metric, conceptualized as representing one or more specific tasks or sets of abilities." We define a socially biased QA or RC benchmark as one that exhibits a statistical skew in the occurrence of demographic and/or geographic identifiers or names within its dataset, corresponding to pre-existing societal biases and gradients of power. Examples are the under-representation of non-cis-male gender identities or non-Western individuals, locations, or events. We would like to address that said skews can be seen as more or less problematic when compared with an assumed ideal distribution, which may differ depending on the purpose of the benchmark or the views of its creator(s) (Shah et al., 2020). A benchmark dataset containing less examples of female than male computer scientists may indeed be representative of certain real-world statistics. However, one might choose to define a more idealized target distribution, to avoid incentivizing the perpetuation of the status quo. Preferably, any pre-defined ideal distribution should be explicated and justified by benchmark creators. Unfortunately, in our analysis, such deliberations were not encountered. Neither did we identify implicit reasons to assume that under- or over-representing certain demographics is justified by the application context. Therefore, we assume uniform ideal distributions in this study. Based on a manual analysis of the 30 most popular QA and RC benchmark papers and a quantitative data analysis of 20 benchmark datasets, our work seeks to answer the following research questions:
@@
    Are the benchmark datasets socially biased? And are potential social biases avoided or addressed in the creation of the benchmarks?
@@
    Are social biases in the datasets reflected in the demographics of the individuals involved in the benchmark creation process?
Our findings are summarized as follows: [^1] The source code can be found here: https://github.com/krangelie/social-bias-qa-benchmarking (RQ1) We identified a lack of transparency regarding demographic details but a general tendency towards Western and, in particular, North American contributors. (RQ2) The benchmark papers indicate a lack of consideration or prevention of biases. Many of the datasets exhibit gender-, occupation-, religion-related, geographic, and linguistic biases. (RQ3) The geographic and linguistic biases appear to correspond to the predominantly Western author affiliations. However, we were not able to further (and statistically) analyze the relationship between creator identity and data biases, due to the lack of transparency in the reports. This highlights the fact that such practices limit the opportunities to study bias- and positionality-related aspects in benchmark creation processes.
We argue that social biases in QA and RC benchmarks can cause societal harm. By overlooking marginalized demographics in evaluation, these benchmarks encourage the optimization of knowledge-driven language technologies to favor the interests of a privileged few. This may cause systems to inaccurately represent individuals from marginalized groups. And it may cause unequal accessibility of relevant knowledge for respective groups, which is a form of epistemic injustice Fricker (2007); Kay et al. (2024); Kraft and Soulier (2024).
@@ 2 Related Works
LLMs reproduce stereotypical associations (Nadeem et al., 2021; Kotek et al., 2023) and achieve different levels of accuracy for examples referring to different social groups in downstream-tasks (Park et al., 2018; Kiritchenko and Mohammad, 2018), such as QA (Parrish et al., 2022; Jin et al., 2024). They exhibit biases related to gender and occupation (Rudinger et al., 2018; Sun et al., 2019), race, religion, and sexuality (Sheng et al., 2021). These biases can lead to representational and allocational harms Barocas et al. (2017); Blodgett et al. (2020). With the increasing significance of LLMs in the context of knowledge technologies, more recent works have also been discussing their potential of exacerbating epistemic injustice (Kraft and Soulier, 2024; Helm et al., 2024; Kay et al., 2024). Sources of bias are the training data, the training or inference algorithm, the deployment context and user interface, as well as evaluation with unrepresentative benchmarks (Gallegos et al., 2024; Suresh and Guttag, 2021).  Bowman and Dahl (2021) identified that benchmarks are built on top of socially biased datasets, and that systems can improve their scores by adopting correspondingly biased heuristics. Raji et al. (2021) criticize that the universality claim of certain AI benchmarks masks their inevitable situatedness and value-ladenness (Haraway, 1988). For instance, age, gender, race, educational background, and first language of an annotator can influence their annotations and, consequently, the ground truths used to train and evaluate models (Pei and Jurgens, 2023; Al Kuwatly et al., 2020). Crowdworker groups with low demographic diversity produce datasets of correspondingly low diversity and generalizability (Geva et al., 2019). Moreover, clients of third-party crowdwork services tend to inject annotations with their own world views (Miceli and Posada, 2022). The situatedness of benchmarks manifests itself in dataset biases, such as in the lack of coverage of "non-Western contexts" (Raji et al., 2021, p.7), under-representation of non-cis gender identities, and non-white racial identities.
@@
Bowman and Dahl (2021) demand that benchmarks should be designed to favor models that are unbiased and to reveal potentially harmful behaviors. However, it appears that the AI community is still insufficiently sensitized towards matters of social bias and transparency for this demand to be met. Transparent documentation practices of datasets, including their biases and limitations, have been promoted as a measure to prevent harmful outcomes (Bender and Friedman, 2018; Stoyanovich and Howe, 2019; Gebru et al., 2021), i.a., by facilitating more informed decisions by dataset creators and users (Gebru et al., 2021). Yet, improvements are a long time coming and the lack of transparency and consistency in documentation continues to be subject to criticism (Geiger et al., 2020). In a structured AI benchmark assessment, Reuel et al. (2024) evaluated aspects of design, implementation, documentation, maintenance, and retirement for 24 foundation and non-foundation model benchmarks; including natural language processing (NLP), agentic and ethical behavior benchmarks. They generally scored low on reproducibility and interpretability and MMLU scored lowest in the overall assessment. Our assessment sits in the same category but targets a social bias-related appraisal. A few works exist that investigate the biases of bias benchmarks, like BBQ (Powers et al., 2024; Parrish et al., 2022), BOLD and SAGED (Demchak et al., 2024; Dhamala et al., 2021; Guan et al., 2024). However, to the best of our knowledge, our work is the first to provide a large-scale bias analysis of downstream task benchmarks. Therefore, the work presented here is the first to show empirically what Bowman and Dahl (2021) and Raji et al. (2021) have warned about on a theoretical level.
@@ 3.1 Benchmark Selection
To identify popular QA and RC benchmarks, we firstly selected all benchmarks including textual data (not excluding multimodal datasets) in the Papers with Code (PwC) corpus of machine learning dataset metadata [^2] https://paperswithcode.com/about, accessed: September 17, 2024. Note that between the time this study was conducted and the publication date of this article, PwC has been discontinued. and ranked them by their citation counts. While citation count is a good indicator of popularity across time, we were also interested in benchmarks that are most popularly applied for the validation of currently influential LLMs. To identify such, we selected the most highly ranked models on the Chatbot Arena LLM Leaderboard (Chiang et al., 2024), [^3] https://lmarena.ai/, accessed: September 18, 2024 as well as the language models with the most likes on HuggingFace. [^4] https://huggingface.co/models?sort=likes, accessed: September 13, 2024 We extracted the top 20 models from both lists and collected all of the 40 related reports, i.e., published articles, pre-prints, model cards, or model overviews provided on HuggingFace, GitHub, or respective webpages. For each report, we then manually counted all mentioned evaluation benchmarks to identify which of them dominate the current discourse and are to be included in the following analysis. Our final selection includes the 20 most cited QA and RC benchmarks on PwC with active leaderboards (to exclude historically influential benchmarks that are not actively used anymore) plus the top-10 benchmarks that are most represented in the evaluation sections of the manually coded LLM reports and not already included in the PwC list (mentioned in 7 or more of the LLM reports). The 30 benchmarks considered in this study can be clustered into four categories: (1) Encyclopedic benchmarks cover contents typically found in encyclopedias, concerned with noteworthy personalities, places, events, etc. Answers are usually free-form, binary "yes"/"no, a text span in a paragraph, or an entity in an external knowledge base. (2) Commonsense benchmarks pose questions about everyday knowledge, e.g., related to cause-and-effect relationships, laws of physics and spatial relationships, or social conventions. Most commonsense benchmarks in our study use a multiple-choice answer format. (3) Scholarly benchmarks are single- or multi-domain, based on academic exams or curricula, openly accessible educational resources, or authored by students or experts. Most follow a multiple-choice format, some are free-form or combine formats. (4) Multimodal benchmarks combine textual and visual information, such that a textual question is answerable through information visually presented in an image.
@@ 3.2 Analysis of Benchmark Papers
Figure 1 gives a schematic overview of our benchmark paper analysis procedure. We followed a content analysis approach similar to Birhane et al. (2022b): we firstly coded all of the benchmark papers, i.e., research articles or introductory pre-prints, [^5] Benchmarks are commonly published on their own, as a byproduct to a technical work, or as a test split to a new corpus.) guided by our research questions. Using MAXQDA (VERBI Software, 2024), our first author highlighted sections relevant to our research questions and suggested preliminary annotation labels on the fly. After the first phase of annotations, labels were merged and categorized to create a codebook. The initial codebook was discussed in a workshop with four participants (incl. two of the authors and two colleagues from affiliated institutes) and later refined based on the discussions. [^6] During the workshop, the codebook was presented as a list of labels with short descriptions and the external participants were asked to annotate two benchmark papers by marking and labeling text spans using this list. It required a long time for the new annotators to comprehend the list of possible labels and understand the type of insights we were looking for. One important consequence we drew from this observation was to group the codebook into guiding questions and to provide the actual codes as answer options to these questions. This helped to accelerate the on-boarding. The final codebook was reformatted and implemented as an online questionnaire via LimeSurvey (LimeSurvey Project Team/ Carsten Schmitz, 2012) for the second wave of annotations. [^7] The full questionnaire is available in our repository. With the final coding schema, all 30 benchmark papers were re-annotated by our first author and one external annotator each. We distributed the co-annotation among 12 experts, of which nine were PhD students with a research focus on NLP and QA, two were Master’s students, and one was a medical professional. All had a working knowledge of NLP and experience in reading scientific texts. All annotators (incl. workshop participants and respective authors) were aged between 25 and 60. They originated from India, Pakistan, China, Germany, and Kazakhstan. All were based in Germany at the time. Roughly one third identified as female. [^8] All annotators (incl. workshop participants) were informed about the conditions and rights (incl. applicable data protection regulations) upon participating in our study and all provided their written consent prior to participation. Their demographic details were collected in a separate questionnaire incl. a separate informed consent form.
@@ 3.3 Analysis of Benchmark Datasets
For the quantitative analysis of social bias within the benchmark datasets, we retrieved external information about entities (people, places, events, etc.) mentioned in the question-answer pairs from Wikidata, [^9] https://www.wikidata.org in particular, gender, occupation, religion, and location-related properties. Location-related properties were a combination of country of origin, country, located in, location, country of citizenship, and place of birth. [^10] We are interested in the general representational tendencies within these various contents and, therefore, group several properties that are indicative of a geographic association. Since, e.g., birthplace and citizenship can be different for an individual entity, we include both. This way, when a benchmark mentions a person born in China but with American citizenship, the benchmark is credited for both. We mostly do not distinguish between human entities and other types of entities, like events and organizations, in our analysis.
@@
Our analysis comprises two different scenarios depicted in Figure 2: Scenario 1: The questions or answers of benchmarks like NaturalQuestions and TriviaQA include entities described in Wikipedia articles and respective identifiers (e.g., article titles or URLs) are provided. Using these identifiers, we queried the Wikipedia API [^11] https://www.mediawiki.org/wiki/API:Main_page to retrieve the corresponding Wikidata QIDs. Using SPARQL, [^12] https://www.w3.org/TR/rdf-sparql-query/ we then retrieved properties of interest for these QIDs directly from the Wikidata knowledge graph, e.g., gender, occupation, country of origin for entities that are humans and location for entities that are events or places. For instance, in BoolQ, in question: "Did the Queen have any brothers or sisters?", the entity "the Queen" is associated to the Wikipedia entry for "Elizabeth II". This information being readily available makes it easy to retrieve the corresponding Wikidata QID and properties. Scenario 2: For benchmarks that are not by default linked to Wikipedia, we applied Entity Linking (EL) against Wikidata using the ReFinED model (Ayoola et al., 2022). [^13] We used the implementation available here: https://github.com/amazon-science/ReFinED (license: Apache 2.0). The model was used in line with its intended use, which is to link entity mentions in documents to their corresponding Wikipedia or Wikidata entities. This allowed us to identify entities mentioned in the questions and retrieve their respective Wikidata identifiers. We then followed the same steps as in the first scenario. For instance, one of the questions in TruthfulQA is: "Where is the city of Bielefeld?" [^14] A correct answer to this question is ”Bielefeld is in Germany” and an expected incorrect answer is ”Bielefeld does not exist”. The entity linker identifies "Bielefeld" as a Wikidata entity with the QID "Q2112", allowing to query properties associated with it from the Wikidata knowledge graph (e.g., its coordinates). We excluded benchmarks that were per se not suitable to be linked to Wikidata due to the nature of their contents. Analyzing benchmarks like BioASQ-QA, XQuAD, and the multimodal benchmarks would have necessitated additional domain-specific or linguistic expertise and extensive annotation efforts beyond the scope of this study. Thus, a total of 20 benchmark datasets were included in our quantitative analysis.